# Hierarchical Diffusion Motion Planning with Task-Conditioned Uncertainty-Aware Priors

### v1

@@ III-B Upper level: key states and timings
The upper level identifies task-relevant key states and their timings, using a lightweight diffusion module parameterized by $\phi$, combined with simple rule-based heuristics as needed. By key states, we mean boundary conditions, waypoints, or states at contact events.
Given a full trajectory $\tau$, let $Y=[y_{1};\ldots;y_{n}]\in\mathbb{R}^{nd}$ be the stack of $n$ key states. Define a binary selection matrix $C\in\{0,1\}^{nd\times Hd}$ that extracts the key states from the full trajectory, so that $Y=C\,\tau$. Note that the selection matrix $C$ is equivalent to the timings for the key states $Y$. The upper level generates these values, and the construction of them depend on the task, as described below.
Maze2D. Key states are evenly spaced temporal waypoints between the start and the goal, so $C$ is fixed by design. In this example, the upper-level diffusion is trained to generate the waypoint set $Y$.
KUKA manipulation. Key states are defined as the joint configuration at the grasp and release timings. The upper-level diffusion is trained to predict the grasp and release timings by generating a binary contact-flag trajectory (1 for contact, 0 for no contact). From this trajectory, we can obtain $C$. The corresponding joint-state set $Y$ is constructed using simple rules (e.g., block-selection heuristics, detailed in Sec. IV) together with inverse kinematics (IK).
The lower level is agnostic to how the high level selects key states. Even if the high level produces imprecise or partially inconsistent $(Y,C)$ with the full trajectories, the lower level treats them as soft observations with uncertainty and proceeds unchanged. Through GP conditioning, this uncertainty enters as observation covariance, so imperfect inputs are down-weighted rather than enforced as hard constraints, and the reverse diffusion remains stable and feasible. For more details, please refer to the subsection below.
III-C Lower level: trajectory generation by denoising
@@
Given the key states $Y\in\mathbb{R}^{nd}$ and the corresponding selection matrix $C\in\{0,1\}^{nd\times Hd}$, we treat $Y$ as soft observations of the stacked trajectory $\tau=[x_{1};\ldots;x_{H}]\in\mathbb{R}^{H\times d}$ with observation covariance $\mathcal{K}_{y}\succeq 0$. The unconditioned GPMP prior (4) induces a GP prior over $\tau$. Conditioning this GP prior to the soft observations $(Y,C,\mathcal{K}_{y})$ yields a task conditioned posterior over the full trajectory,
$\tau(t)\sim\mathcal{N}\big(\mu(t),\,\mathcal{K}(t,t^{\prime})\big),$ (8)
with a closed form mean and covariance [10]
@@
Small variances in $\mathcal{K}_{y}$ enforce tight adherence to key states (e.g., fixed initial or goal states), while larger variances allow flexibility (e.g., soft waypoints). Intuitively, $Y$ anchors the trajectory near important states and $\mathcal{K}_{y}$ encodes confidence in those anchors, without imposing hard constraints.
The conditioned GP (8) determines the non-isotropic and time-correlated Gaussian noise used in our diffusion model; at each step, the injected noise is drawn with a mean and covariance shaped by linear time varying dynamics and the key state and timing information. In particular, the lower level initializes reverse diffusion from $\mathcal{N}(\mu,\mathcal{K})$ and denoises under the general forward and posterior relations introduced in Sec. II-A. Recall that the one-step forward posterior for our biased and non-isotropic setting has the closed form
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\tilde{\mu}_{i}(\tau^{i},\tau^{0},\mu),\,\tilde{\beta}_{i}\,\mathcal{K}\big),$ (11)
where $\tilde{\mu}_{i}$ and $\tilde{\beta}_{i}$ generalize the DDPM posterior coefficients $\hat{\mu}_{i}$ and $\hat{\beta}_{i}$. For learning we approximate the reverse kernel
@@
where cond may include $Y$, $C$, initial states, and other context, and we take $\Sigma^{i}$ to match the scaled structure $\tilde{\beta}_{i}\,\mathcal{K}$. Following [9] and Sec. II-A, minimizing the KL divergence between $p_{\theta}$ and the true forward posterior reduces to a weighted regression toward $\tilde{\mu}_{i}$. Using the Mahalanobis norm $\|v\|^{2}_{A}:=v^{\top}A\,v$, the training objective becomes
@@
up to a scalar schedule dependent factor. Compared to the isotropic MSE, (12) naturally weights errors by the inverse of the task conditioned covariance, encouraging denoising steps that stay consistent with the GP induced temporal coupling and the confidence encoded in $\mathcal{K}_{y}$.
III-D Pipeline Summary
The complete training process is shown in Alg. 1 and Alg. 2. For the upper level, we first obtain the ground truth $(C,Y)$ and choose the supervision targets, which are task dependent We then update the network parameters of the upper-level diffusion with the standard diffusion loss on the chosen target(s) (Alg. 1). For the lower level, we use the ground truth $(C,Y)$ to instantiate the task specific prior by computing $(\mu,\mathcal{K})$ via (9)–(10). We then update the network parameters of the lower-level diffusion with the Mahalanobis MSE (Alg. 2).
The complete inference process is shown in Alg. 3. At sampling time, the upper level outputs $(C,Y)$, from which we construct $(\mu,\mathcal{K})$. At the lower level, we use $(\mu,\mathcal{K})$, initialize $\tau^{N}\sim\mathcal{N}(\mu,\mathcal{K})$, and run the reverse chain using (7) with the learned denoiser until a fully denoized trajectory is obtained.
@@ III-D Pipeline Summary
Algorithm 1 Upper-level training (for key states and/or timings)
1:Inputs: dataset $\mathcal{D}$ of trajectories $\tau$
2:Initialize upper-level diffusion module parameters $\phi$
@@
4:  Sample $\tau\sim\mathcal{D}$
5:  Extract ground-truth supervision targets $(C,Y)$ $\triangleright$ depending on application, train on $C$, $Y$, or both
6:  Update $\phi$ using the standard diffusion loss w.r.t. the chosen target(s)
@@
4:  Sample $\tau\sim\mathcal{D}$, set $\tau^{0}=\tau$, and extract ground truth $(C,Y)$
5:  Form $(\mu,\mathcal{K})$ via (9)–(10) (usually extract from cached pre-computed ones)
6:  Sample $i\sim\text{Unif}\{1,\dots,N\}$
7:  $\tau^{i}\!\sim\!\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{i}}\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\mu,\;(1-\bar{\alpha}_{i})\mathcal{K}\right)$
8:  Gradient descent on $\nabla_{\theta}\left\|\tilde{\mu}_{i}(\tau^{i},\tau^{0})-\mu_{\theta}(\tau^{i},i)\right\|_{\mathcal{K}^{-1}}$
9:end while
Algorithm 3 Test-time Planning: Hierarchical Diffusion
@@
3:Prior build: $(\mu,\mathcal{K})\leftarrow\text{GP-Post}(\tilde{\mu},\widetilde{\mathcal{K}},C,Y,\mathcal{K}_{y})$ via (9)–(10)
@@
5:for $i=N,\dots,1$ do
6:  $\tau^{i-1}\sim\mathcal{N}(\mu_{\theta}(\tau^{i},i),\tilde{\beta}_{i}\mathcal{K})$
7:end for
8:return $\tau^{0}$
@@
Motion Planning Diffusion (MPD) [1] applies a GP as a guidance cost only during reverse sampling, while the forward corruption remains isotropic. We instead parameterize the noise with a GP: conditioning on key states sets the mean $\mu$ and covariance $\mathcal{K}$ that govern both the forward and posterior distributions. This yields three benefits. First, structure is present at every step of training and sampling, so denoising follows a Mahalanobis geometry that narrows the search to trajectories consistent with temporal coupling and task constraints. Second, uncertainty in key states enters as an observation covariance $\mathcal{K}_{y}$, which downweights imperfect cues without hard constraints. Third, feasibility comes from the generative model rather than from external costs, reducing sensitivity to guidance weights and improving stability over long horizons.
@@
Figure 4: Maze2D goal reaching. Top-Left: task map with start (blue) and goal (red). Top-right: baseline (isotropic Gaussian noise) samples. Bottom: our method. The small inset on the left shows the waypoints predicted by the upper level (see Fig. 2). Although waypoints provide a rough guide for the global trajectory, they can be imperfect or even infeasible (with collisions), so hard-constraining the lower level using these waypoints is brittle. Thus, we treat them as noisy observations and encode their uncertainty in the GPMP prior with the observation covariance $\mathcal{K}_{y}$, which guides denoising to yield smooth, collision-free trajectories aligned with the task.
@@ IV Experiments
We adopt Diffuser [2] as the baseline and evaluate the proposed structured-noise formulation on two domains: (i) Two-dimensional maze navigation (Maze2D) with initial and goal conditioning, and (ii) KUKA block stacking with grasp and release events represented with contact flag trajectory for each block. We report goal success, dynamic feasibility, and constraint satisfaction.
@@ IV-A Experimental Setup
IV-A1 Comparison models
We compare four planners that share the same network backbone and differ only in how they use key states and in the structure of the diffusion noise:
  • Isotropic (without key states). A standard DDPM with isotropic Gaussian corruption and no high-level key states (the flat Diffuser baseline).
  • Isotropic (with key states as conditioning). A standard DDPM at the lower level that receives the high-level key states as conditioning inputs, but still uses isotropic Gaussian corruption at the low level.
  • GPMP prior (without key states). A single-level diffuser with non-isotropic Gaussian noise shaped by the GPMP prior, but without key states. This highlights that relying only on a GP prior for dynamic feasibility can harm trajectory generation by causing oversmoothing.
  • GPMP prior (with key states). Our method: a two-level hierarchy in which the upper level produces task-centric key states and timings, and the lower level denoises under the non-isotropic Gaussian from GPMP.
Note that the choice of noise prior applies only to the lower-level diffusion. Whenever a hierarchy is used, the upper-level diffusion remains a standard DDPM with isotropic Gaussian noise.
IV-A2 Neural network architecture
Both levels use a UNet architecture as in [2]. For the lower level, the network is conditioned on the key states and their timings produced by the upper level. Apart from this conditioning and the choice of noise model (isotropic versus GP-conditioned), the architecture and training schedule are identical across all models so that any performance differences are induced by the effect of the noise structure.
IV-A3 Maze2D goal reaching
@@
Each task specifies a start and a goal state on a fixed maze map. The upper-level diffusion produces temporal waypoints $Y$ at fixed times $C$, which are treated as soft observations with associated uncertainty. These waypoints instantiate the task-conditioned GP prior, from which the lower level generates full trajectories by reverse diffusion.
IV-A4 KUKA arm block stacking
@@
The task is grasp-and-place with a 7-DoF KUKA arm. The dataset provides 7-DOF joint position trajectories and a binary contact-flag trajectory (one bit per time step) for each block, while block poses are determined deterministically from kinematics and contact.
@@
As explained in Sec.III, the key states are grasp and release joint positions. Given a ground-truth trajectory $\tau^{0}$ and its contact flags, we define the grasp configuration as the joint state at the first $0\!\to\!1$ transition and the release configuration as the joint state at the first $1\!\to\!0$ transition. The stack $Y$ consists of the initial joint state together with these grasp and release joint states, while $C$ is taken to be the binary contact-flag trajectory over the horizon. The upper-level diffusion is trained to generate $C$ conditioned on $Y$.
@@
At test time, we input the grasp and stack targets with the following heuristics: (1) if a block is currently grasped, use it as the grasp target; (2) choose the release target as the block with the largest vertical position, excluding any grasped block; (3) if no block is grasped, choose the block closest to the initial end-effector position; (4) if the chosen grasp–release pair is already stacked ($|dx|<\mathrm{tol}$, $|dy|<\mathrm{tol}$), select a different grasp block. After the block selection, we compute the grasp and release joint states by IK to form $Y$. Then, the upper-level diffusion predicts the contact-flag trajectory $C$; grasp and release timings are extracted from the first $0\!\to\!1$ and $1\!\to\!0$ transitions, respectively.
IV-B Results
IV-B1 Success Rate
In Maze2D, we define a trajectory as successful if two conditions are met: (i) the final state lies within a predefined tolerance radius around the goal, and (ii) the entire trajectory remains collision-free with respect to the maze map. For each planning model, results are averaged over 100 randomly sampled start–goal pairs.
@@ IV-B1 Success Rate
In KUKA block stacking, success requires both a valid grasp and a valid placement within tolerance. We declare a grasp successful when the end-effector is within a spatial threshold of the target block at the first contact onset (flag transition $0\!\to\!1$). A placement is declared successful when, at the release time (flag transition $1\!\to\!0$), the end-effector is within tolerance of the designated stack location. Results are averaged over 100 randomized initial joint configurations and block layouts.
We evaluate a $2\times 2$ grid of diffusion planners that vary along two axes: (i) isotropic Gaussian noise versus the GPMP prior (structured, task-aware mean $\mu$ and covariance $\mathcal{K}$), and (ii) without versus with high-level key states. All models share the same UNet backbone, are trained for the same number of optimization steps, and are verified to have comparable convergence in training loss.
Tables I and II report the success rates. Across both tasks, the best performance is achieved by combining the upper-level key-state generator with the GPMP prior. Intuitively, upper-level key states and timings supply task semantics (e.g., guidance through turns in a maze or contact phases in manipulation), while the GPMP prior encodes smoothness and soft constraint satisfaction. Together, these two components concentrate probability mass of the reverse process near dynamically feasible and task-consistent manifolds, thereby reducing the burden on the denoiser.
Interestingly, the GPMP prior without informative key states underperforms even the isotropic baseline. This occurs because, although the GPMP prior captures temporal correlations and promotes smoothness, it can oversmooth when applied uniformly across the entire horizon. Realistic trajectories often involve non-stationarities: sharp turns in navigation and abrupt grasp/release events in manipulation. Without explicit key states to anchor the dynamics, the GP prior may bias trajectories away from the task-critical states, leading to degraded success. Effective planning, therefore, requires a balance: the hierarchical key-state generator anchors task-critical events, while the prior provides tighter variance around the key states ensuring they are roughly respected, while maintaining smoothness and temporal consistency. The combination of both yields robust, dynamically feasible, and task-aligned plans.
Maze2D (100 samples) Success
Isotropic (without key states) 36/100
Isotropic (with key states as conditioning) 42/100
GPMP Prior (without key states) 14/100
GPMP Prior (with key states) (our model) 75/100
TABLE I: Maze2D success rates (number of collision-free successful trajectories out of 100) after 8080k training steps. Model names match the definitions in Sec. IV.
KUKA stacking (100 samples) Success
Isotropic (without key states) 11/100
Isotropic (with key states as conditioning) 12/100
GPMP Prior (without key states) 8/100
GPMP Prior (with key states) (our model) 71/100
TABLE II: KUKA stacking success rates (number of successful grasps and placements out of 100) after 3030k training steps. Model names match the definitions in Sec. IV.
@@
Figure 6: Maze2D velocity consistency. Left: isotropic Gaussian prior; Right: our model. Blue: generated velocities; Orange: numerical derivatives of generated positions. Closer alignment indicates better dynamic feasibility, achieved with the GPMP prior.
Maze2D (100 samples) Mean Absolute Error
Isotropic (without key states) 0.22
Isotropic (with key states as conditioning) 0.21
GPMP Prior (without key states) 0.08
GPMP Prior (with key states) (our model) 0.08
@@
TABLE III: Maze2D velocity consistency: mean absolute error (MAE) between generated velocities and finite-difference derivatives of positions over 100 trajectories. With GPMP noise, MAE remains low regardless of key states, since the LTV-GP prior enforces natural position–velocity alignment.
IV-B2 Dynamic Feasibility & Smoothness
We next evaluate whether the generated trajectories satisfy dynamic feasibility and exhibit smoothness.
Maze2D.
The Diffuser baseline provides position, velocity, and action effort sequences. While the success rate is computed only from position plans, our method is trained with a GPMP-inspired prior that couples position and velocity, thereby encoding feasibility. To assess this property, we compare the generated velocities against numerical derivatives of the generated positions (finite difference over the time step, Fig. 6). We also report the mean absolute error (MAE) along the horizon (Table III). We additionally present representative trajectories (Fig. 4) along with their denoising processes (Fig.  3), providing qualitative insight into the behavior of the models. Qualitatively, our method generates smooth and feasible motions, whereas the isotropic baseline often produces jagged position traces with discontinuities inside the trajectories (caused by hard enforcement of start and goal conditions), resulting in unrealistic and infeasible plans .
KUKA block stacking.
The open-source baseline does not provide velocities or actions. Thus, feasibility cannot be directly evaluated in this domain, but smoothness can still be assessed by measuring the proximity between consecutive joint states along the trajectory. Visualizations confirm that trajectories generated with our structured prior are smoother and more coherent compared to the baseline (Fig. 5).
V Discussion and Conclusion
Embedding GPMP structure in the noise yields temporally consistent, constraint-aware trajectories, improving success under sparse or imperfect high-level guidance. Our framework is task-centric: the upper level extracts task-relevant key states/timings, and the lower-level diffusion uses the GP-conditioned $(\mu,\mathcal{K})$ to bias sampling toward feasible solutions. We validate the approach through experiments on two robotic domains. In Maze2D, the prior concentrates probability mass near start/goal and along smooth waypoint-guided paths; in KUKA stacking, predicted contact timing shapes the kernel so denoising honors grasp/release phases. Although we proposed a lightweight diffusion model for the upper level, the lower level is agnostic to how the key states are obtained—be it rule-based, model-based, or learned. It always treats $(Y,C)$ as soft observations with uncertainty; GP conditioning down-weights imperfect inputs rather than enforcing them as hard constraints, leaving room to pair this framework with alternative task specifiers or high-level motion planners.
Moreover, our method does not require handcrafted reward terms or auxiliary objectives, and no advanced neural architectures are used beyond conditioning on upper-level-generated features. Thus improvements in sample quality and task success arise from where the structure lives—in the noise—rather than from heavier networks or guidance heuristics. This shows our method’s efficacy even with simple mechanics.

### v2

@@ III-A Upper level: key states and timings
The upper level identifies task-relevant key states and their timings, with simple rule-based heuristics as needed. By key states, we mean boundary conditions, waypoints, or states at contact events.
Given a full trajectory $\tau\in\mathbb{R}^{Hd}$, let $Y=[y_{1};\ldots;y_{n}]\in\mathbb{R}^{nd}$ be the stack of $n$ key states. Define a binary selection matrix $C\in\{0,1\}^{nd\times Hd}$ that extracts the key states from the full trajectory, so that $Y=C\ \tau$. Note that the selection matrix $C$ directly encodes the timings of the key states $Y$. The way we construct supervision target for $(C,Y)$ is task-dependent, and is described in Sec. IV. The complete training process is shown in Alg. 1. We first obtain the ground truth $(C,Y)$ and choose the supervision targets, which are task dependent. We then update the network parameters of the upper-level diffusion with the standard diffusion loss on the chosen target(s).
III-B GPMP prior construction
@@
Given the key states $Y\in\mathbb{R}^{nd}$ and the corresponding selection matrix $C\in\{0,1\}^{nd\times Hd}$, we treat $Y$ as soft observations of the stacked trajectory $\tau=[x_{1};\ldots;x_{H}]\in\mathbb{R}^{H\cdot d}$ with observation covariance $\mathcal{K}_{y}\succeq 0$. The unconditioned GPMP prior (4) induces a GP prior over $\tau$ with mean $\tilde{\xi}=[\tilde{\xi}_{1};...;\tilde{\xi}_{H}]\in\mathbb{R}^{H\cdot d}$ and covariance matrix $\tilde{\mathcal{K}}=[\tilde{\mathcal{K}}(t_{i},t_{j})]_{1\leq i,j\leq H}$, where subscript $i$ represents discretized time step $i$. Conditioning this GP prior to the soft observations $(Y,C,\mathcal{K}_{y})$ yields a task conditioned posterior over the full trajectory with a closed form mean and covariance [18]
@@
This posterior $\mathcal{N}(\xi,\mathcal{K})$ serves as the task-conditioned prior for the lower-level reverse diffusion. Small variances in $\mathcal{K}_{y}$ enforce tight adherence to key states (e.g., fixed initial or goal states), while larger variances allow flexibility (e.g., soft waypoints). Intuitively, $Y$ anchors the trajectory near important states and $\mathcal{K}_{y}$ encodes confidence in those anchors, without imposing hard constraints.
III-C Lower level: trajectory generation by denoising
The conditioned prior $\mathcal{N}(\xi,\mathcal{K})$ determines the non-isotropic and time-correlated Gaussian noise used in our diffusion model; at each step, the injected noise is drawn with a mean and covariance shaped by linear time varying dynamics and the key state and timing information. In particular, the lower level initializes reverse diffusion from $\mathcal{N}(\xi,\mathcal{K})$ and denoises under the general forward and posterior relations introduced in (6) – (7). For learning, we approximate the reverse posterior (7) with
@@ III-C Lower level: trajectory generation by denoising
where cond may include $Y$, $C$, initial states, and other context, and we take $\Sigma^{i}$ to match the scaled structure $\tilde{\beta}_{i}\,\mathcal{K}$. Following [10] and Sec. II-A, minimizing the KL divergence between $p_{\theta}$ and the true reverse posterior reduces to a weighted regression toward $\tilde{\mu}_{i}$. Using the Mahalanobis norm $\|v\|^{2}_{A}:=v^{\top}A\,v$, the training objective becomes
@@
up to a scalar schedule dependent factor. This weights errors by the inverse of the task conditioned covariance, encouraging denoising steps that stay consistent with the GP induced temporal coupling and the confidence encoded in $\mathcal{K}_{y}$.
During lower-level training, we use the ground-truth key states and timings $(C,Y)$, and perturb them to $(C+dC_{w},\,Y+dY_{w})$ to account for imperfections in the upper-level predictions. The perturbed values are used to instantiate the task-specific prior $(\xi,\mathcal{K})$ via (8)–(9). We then update the lower-level diffusion parameters using the Mahalanobis MSE loss (10), as summarized in Alg. 2. The complete inference process, including both the upper and lower levels, is summarized in Alg. 3.
@@
Algorithm 1 Upper-level training for key-state and timing prediction
1:Inputs: dataset $\mathcal{D}$ of trajectories $\tau^{0}$
2:Initialize upper-level diffusion parameters $\phi$
@@
4:  Sample trajectory $\tau^{0}\sim\mathcal{D}$
5:  Extract supervision targets $(C,Y)$ from $\tau^{0}$ $\triangleright$ $C$: timing indices, $Y$: key states
6:  Update $\phi$ using the standard diffusion loss to predict $C$, $Y$, or both
@@
4:  Sample $\tau^{0}\sim\mathcal{D}$ and extract $(C,Y)$ from $\tau^{0}$
5:  $(C,Y)\leftarrow(C+dC_{\omega},\;Y+dY_{\omega})$ $\triangleright$ $dC_{\omega},dY_{\omega}$ are random perturbations
6:  $(\xi,\mathcal{K})\leftarrow\text{GP-Post}(\tilde{\xi},\widetilde{\mathcal{K}},C,Y,\mathcal{K}_{y})$ via (8)–(9)
7:  Sample $i\sim\text{Unif}\{1,\dots,N\}$
8:  $\tau^{i}\!\sim\!\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{i}}\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\xi,\;(1-\bar{\alpha}_{i})\mathcal{K}\right)$
9:  Gradient descent $\nabla_{\theta}\left\|\tilde{\mu}_{i}(\tau^{i},\tau^{0},\xi)-\mu_{\theta}(\tau^{i},i)\right\|_{\mathcal{K}^{-1}}^{2}$
10:end while
Algorithm 3 Test-time Planning (Inference)
@@
4:Lower Level:
[…]
6:for $i=N,\dots,1$ do
7:  $\tau^{i-1}\sim\mathcal{N}(\mu_{\theta}(\tau^{i},i),\tilde{\beta}_{i}\mathcal{K})$
8:end for
9:return $\tau^{0}$
@@
Motion Planning Diffusion (MPD) [4] applies a GP as a guidance cost only during reverse sampling, while the forward corruption remains isotropic. We instead parameterize the noise with a GP: conditioning on key states sets the mean $\xi$ and covariance $\mathcal{K}$ that govern both the forward and posterior distributions.
@@
Figure 3: Maze2D goal reaching. Top-Left: task map with start (blue) and goal (red). Top-right: baseline (Diffuser) samples. Bottom: our method. The small inset on the left shows the waypoints predicted by the upper level, which might be imperfect. Thus, instead of hard constraints, we model them as noisy observations.
@@
Figure 5: Average roughness (AR) and success rate over training for Maze2D (left) and KUKA (right). Metric calculated as average of 10 samples. The dashed line shows the dataset-average AR as a reference for reasonable smoothness.
@@ IV-A Experimental Setup
We evaluate the proposed structured-noise formulation on two domains. We describe the benchmark tasks and how task-specific key states and timings are defined.
@@
(i) Maze2D goal reaching. Each task specifies a start and a goal state on a fixed maze. In this example, key states are evenly spaced temporal waypoints between the start and the goal, so $C$ is fixed by design since the waypoint timings—and thus the selection indices—are predetermined. The upper-level diffusion is trained to generate the waypoint set $Y$ at fixed times $C$. These waypoints instantiate the task-conditioned GP prior, from which the lower level generates full trajectories by reverse diffusion.
@@
(ii) KUKA arm block stacking. The task is grasp-and-place with a 7-DoF KUKA arm. The dataset provides joint-position trajectories together with a binary contact-flag trajectory (one bit per time step, where 1 indicates contact) for each block; block poses are determined deterministically from kinematics and contact.
@@
Key states are defined as the joint configurations at grasp and release events. Accordingly, the upper-level diffusion is trained to predict the contact-flag trajectory, from which grasp and release timings are extracted. Given a ground-truth trajectory $\tau^{0}$ and its contact flags, the grasp configuration is defined as the joint state at the first $0\!\to\!1$ transition, and the release configuration as the joint state at the first $1\!\to\!0$ transition. The key-state stack $Y$ consists of the initial joint state together with these grasp and release configurations, while the selection matrix $C$ is given by the contact-flag trajectory over the horizon.
@@
At test time, grasp and release block targets are selected using simple heuristics: (1) if a block is currently grasped, it is used as the grasp target; (2) the release target is chosen as the block with the largest vertical position, excluding any grasped block; (3) if no block is grasped, the block closest to the initial end-effector position is selected; (4) if the chosen grasp–release pair is already stacked, a different grasp block is selected. After block target selection, the corresponding grasp and release joint states are computed via inverse kinematics (IK) to form $Y$, and the upper-level diffusion predicts the contact-flag trajectory $C$.
IV-B Baselines and Ablations
We compare three planners in addition to ours that share the same UNet backbone and training schedule, differing only in how task structure is incorporated. All models are trained for the same number of optimization steps (100k steps with batch size 32), and we verified that their training losses converged, ensuring that performance differences arise from the choice of noise structure and task representation.
The detailed design of each planner is as follows:
(i) Diffuser. The baseline Diffuser [12], which uses isotropic Gaussian noise and conditions only on task inputs such as initial and goal states.
(ii) Diffuser + KeyCond. A hierarchical variant of Diffuser where the lower-level diffusion is conditioned on key states and their timings produced by the upper level, while still using isotropic Gaussian noise. This isolates the effect of neural task conditioning without structured noise.
(iii) EDMP. A single-level diffuser with isotropic Gaussian noise, guided at inference time by model-based cost functions. Following the general framework of [23], we design task-specific cost functions for each benchmark. For Maze2D, the cost penalizes collisions by summing penetration depths of each state with the maze obstacles, together with penetration depths at the midpoints between adjacent states to approximate swept-volume collisions. This implementation requires access to the maze geometry at test time. For KUKA block stacking, since [23] does not specify a cost for grasp-and-place tasks, we design a simple heuristic cost based on the minimum distance between the end-effector and target blocks, rewarding configurations with small distnace.
(iv) Hierarchical GPMP (Ours). Our two-level approach, where the upper level predicts key states and timings, and the lower level denoises under a GPMP-conditioned prior.
Note that the choice of noise prior applies only to the lower-level diffusion. Whenever a hierarchical model is used, the upper-level diffusion always employs a standard DDPM with isotropic Gaussian noise.
IV-C Results and Analysis
@@
Task encoding with GPMP conditioning enables high success rates. In Maze2D, a trajectory is considered successful if (i) the final state lies within a predefined tolerance radius of the goal and (ii) the trajectory remains collision-free with respect to the maze map. Results are averaged over 100 randomly sampled start–goal pairs. In KUKA block stacking, success requires both a valid grasp and placement within tolerance. A grasp is successful if the end-effector is within a spatial threshold of the target block at the first contact onset (flag transition $0\!\to\!1$). Placement is successful if, at release time, the end-effector lies within tolerance of the designated stacking location. Results are averaged over 100 randomized initial joint configurations and block layouts.
Table I shows that our method achieves the best success rates across both tasks. This improvement can be attributed to the structured prior, which concentrates probability mass near task-critical trajectories through soft conditioning in (8)–(9).
The results of Diffuser + KeyCond further indicate that the primary benefit arises from the structured noise model rather than from neural conditioning alone. We encode $(Y,C)$ with a lightweight MLP and use the resulting features to condition the diffusion model for both Diffuser + KeyCond and our method, enabling a controlled comparison that isolates the effect of the noise structure.
Method Maze2D KUKA Stacking
Diffuser 48/100 31/100
Diffuser + KeyCond 49/100 28/100
EDMP 56/100 34/100
Hierarchical GPMP (Ours) 81/100 70/100
TABLE I: Task success rates over 100 samples.
@@
Figure 6: Maze2D velocity consistency. Left: isotropic Gaussian prior; Right: our model. Blue: generated velocities; Orange: numerical derivatives of generated positions. Closer alignment indicates better dynamic feasibility.
Maze2D (100 samples) Velocity MAE
Diffuser 0.2187
Diffuser + KeyCond 0.2064
EDMP 0.2012
Hierarchical GPMP (Ours) 0.0767
@@
TABLE II: Maze2D velocity consistency: mean absolute error (MAE) between generated velocities and finite-difference derivatives of positions over 100 trajectories.
LTV-GP-based covariance produces smoother trajectories. We evaluate the smoothness of the generated trajectories using both qualitative and quantitative measures. For quantitative evaluation, we adopt the Average Roughness (AR) metric [23], defined as the average of $\|x_{t+1}-x_{t}\|$ along the trajectory. Lower AR corresponds to smoother trajectories; however, excessively low values (e.g., AR = 0, corresponding to no movement) are not desirable. We therefore compare the AR of generated trajectories against the ground-truth AR computed from the dataset.(See Fig. 5). We additionally present representative trajectories (Fig. 3, Fig. 4) and their denoising processes (Fig. 2) for qualitative comparison.
LTV-GP covariance captures inter-state dependencies. We evaluate whether generated trajectories preserve dependencies between state dimensions. In the Maze2D benchmark, the Diffuser baseline outputs position, velocity, and action-effort sequences. Our method instead employs a GPMP-inspired prior that explicitly couples position and velocity through the LTV-GP covariance, encouraging dynamically feasible trajectories. To assess this property, we compare predicted velocities with finite-difference derivatives of positions (Fig. 6) and report the mean absolute error (MAE) over the trajectory horizon (Table II).
Overall, our method produces trajectories with stronger position–velocity consistency, whereas the baseline often exhibits jagged position traces and inconsistent velocity profiles—especially near hard start and goal constraints—leading to dynamically infeasible behavior.
Encoding task information in the prior improves training efficiency. The structured prior preserves task-related information during the corruption process, leading to more efficient training. We report average roughness (AR) and success rate (SR) over training steps in Fig. 5. While the vanilla Diffuser gradually produces smoother trajectories as training progresses, our method achieves dataset-level smoothness from early stages due to the structured prior, whereas the baseline requires substantially more training to reach comparable AR values.
A similar trend is observed for SR: our method attains high SR at early training stages. Notably, desirable trajectory properties such as smoothness and task success do not necessarily converge with the diffusion training loss. Thus, relying on training loss alone makes it difficult to identify a desirable operating point. Although longer training may eventually improve these metrics, diffusion models are computationally expensive, making such prolonged training inefficient.
Separating timing prediction enables modeling phase-dependent task structure. In many robotics tasks, desired behavior depends not only on the time index but also on the task phase (e.g., grasp, transport, and release in the KUKA example). Encoding such phase-dependent behavior through cost guidance is challenging. While collision-avoidance costs can evaluate each state independently (e.g., checking whether a state is in collision), tasks that require interactions at precise moments create strong coupling between timing variables (e.g., the contact-flag trajectory) and robot joint states. Moreover, cost-guided diffusion must evaluate partially denoised trajectories during sampling, where interaction times can shift significantly under noise. These factors make such non-stationary task structures difficult to represent with a differentiable cost.
Our framework addresses this by assigning timing prediction to the upper level while the lower level generates trajectories conditioned on these predictions with uncertainty. This separation enables more effective modeling of phase-dependent behaviors and contributes to the improved performance in Table I.
V Conclusion
In this paper, we propose a hierarchical diffusion planner that embeds task and motion structure directly in the noise model. We validate the approach through experiments on two robotic domains, where our method demonstrates improved task success and trajectory quality, such as smoother trajectories and better inter-state coupling consistency.
Although we use a lightweight diffusion model at the upper level, the lower level is agnostic to how the key states are obtained—be it rule-based, model-based, or learned. Moreover, our method does not require handcrafted reward terms or auxiliary objectives for guidance, and it does not rely on advanced neural architectures beyond conditioning on upper-level-generated features.
Future work includes applying the framework to more complex robotic systems (e.g., humanoids) and incorporating more advanced task specifiers, such as VLM-based modules, at the upper level.